# Probing Deep into Temporal Profile Makes the Infrared Small Target Detector Much Better

### v1

@@ 3.2.1 Attribution approach for IRST detection networks
[…]
@@
Based on integrated gradients, we developed the first prediction attribution analysis tool for IRST detection networks. Unlike classification networks, segmentation-based IRST detection networks generate the prediction of each spatial pixel in the input $\mathbf{S}\in\mathbb{R}^{T\times H\times W}$, and can be represented as $D:\mathbb{R}^{T\times H\times W}\longrightarrow\mathbb{R}^{H\times W}$. $T$ denotes the temporal length of the input, while $H\times W$ denotes the spatial size of the input. To interpret $D$, we attribute the global prediction of a single target. Specifically, we need a converter to quantify the prediction of a target, $C:\mathbb{R}^{H\times W}\longrightarrow\mathbb{R}$. In this work, we achieved the converter by summing the predictive scores of all pixels belonging to one target. That is,
@@
[…]
where $\mathbf{S^{\prime}}\in\mathbb{R}^{T\times H\times W}$. Since the targets always present as high-frequency components in the input, the baseline should avoid containing high-frequency components. Meantime, arbitrary fluctuation in the background can affect the prediction of the target. Therefore, the constant tensor is suitable as a baseline line for IRST detection, and the constant should be lower than the gray average of the target since integrated gradients fail to highlight black targets [83]. The path used in this work is a straight line, i.e.,
@@ 3.3 Temporal probe mechanism
Figure 8: Visualization of the structure of TPro. Firstly, temporal probes are employed to pull out complete temporal features of a single pixel from the split input features. Then, multiple learnable SCorMs are applied to the gotten features to extract temporal correlation features. This process is performed pixel-wise across all split input features. Finally, all temporal correlation features are concatenated and fused to generate output features. Notably, the entire pipeline strictly excludes any space-dimension computations.
Inspired by the above research, we remodel the IRST detection task as a one-dimensional (i.e., time-dimension) signal anomaly detection task. The target signals are treated as temporal anomalies. As previously analyzed, it is important for the networks to extract long-term temporal correlation information for this task. Ordinary 3D convolution operations focus only on short-term features and require stacking multiple convolution layers to achieve long temporal coverage. However, the temporal receptive field gradually expands as the number of stacked layers increases, and thus a amount of feature extraction computation is still based on short-term features. Moreover, multiple 3D convolutional layers bring substantial computational overhead for large-scale feature tensors. Therefore, in this part, we propose a temporal probe mechanism (TPro) to efficiently extract long-term temporal correlation features and to probe whether the temporal features of a pixel exist anomaly. The structure of TPro is shown in Fig. 8.
@@
As shown in Fig. 8, the input basic features are first divided into $m$ features along the channel dimension, i.e., $F^{b}_{i}\in\mathbb{R}^{\frac{C}{m}\times T\times H\times W}$ ($i=1,2,...,m$). To perform computations on the long-term temporal features of a pixel, temporal probes are utilized to individually extract the complete temporal features of each pixel, i.e., $F^{b}_{i}(x,y)\in\mathbb{R}^{\frac{C}{m}\times T\times 1\times 1}$ ($x\in[1,W],y\in[1,H]$). Subsequently, the correlation matrix is applied to the extracted features for the temporal correlation features $F^{cor}_{i}(x,y)\in\mathbb{R}^{\frac{C}{m}\times T\times 1\times 1}$. The process can be described as
@@
where $W_{i}^{T\times T}$ represents the $i^{th}$ correlation matrix. After repeating the above process, the temporal correlation features (i.e., $F^{cor}_{i}$) of all pixels are obtained. To fuse temporal correlation information from different channels and correlation matrixes, a pointwise convolution [105] (P-Conv, i.e., $1\times 1\times 1$ convolution) layers $\phi(\cdot)$ is applied to these features (i.e., $F^{cor}_{1},F^{cor}_{2},...,F^{cor}_{m}$) after concatenation along the channel dimension. This process can be described as
@@ 3.4 Deep temporal probe network for IRST detection
Specifically, tens of frames $\mathbf{S}$ are first fed to a temporal convolution (T-Conv) layer [^4] Each convolution layer consists of a convolution operation, batch normalization, and ReLU activation. with a kernel $\mathbf{W_{T}}\in\mathbb{R}^{l\times 1\times 1}$ ($l=5$). Then, the features are sent to a few identical temporal difference residual blocks (TD-ResBlock) to roughly extract the basic temporal features $F^{b}$. Each TD-ResBlock contains a temporal difference convolution (TD-Conv) layer $\psi(\cdot)$ with a $3\times 1\times 1$ kernel $\mathbf{W_{TD}}$ and a temporal dilation of 2, a T-Conv layer ($l=3$), a P-Conv layer and the skip connection. TD-Conv can maintain and emphasize the temporal saliency of targets roughly by applying convolution to temporal difference features (i.e., $f_{t}-f_{t-1}$), which can be described as
@@ 4.1.1 Datasets
TABLE I: Parameters related to noise of the datasets. The standard deviation of noise ($\sigma_{n}$), the distribution range of the multiplicative noise ($\sigma_{g}$), the standard deviation of the additive noise ($\sigma_{o}$), the pattern of non-uniformity, and the average SNR value of the whole dataset.
Datasets $\sigma_{n}$ $\sigma_{g}$ $\sigma_{o}$ non-uniformity SNR
NUDT-MIRSDT 0.1 0 0 none 6.40
NUDT-MIRSDT-HiNo 8.0 0.15 1.3 pixel-wise 2.66
@@ 4.1.2 Metrics
We followed [7] to use the probability of detection ($P_{d}$), false alarm rate ($F_{a}$), and area under curve (AUC) as quantitative metrics for the evaluation of IRST detection performance. $P_{d}$ measures the ability of a method to detect targets. $F_{a}$ measures the ability of a method to correctly identify the background, such as clutter. AUC is the area value under the receiver operating characteristics (ROC) curve [109, 110], reflecting the detection performance under different thresholds. Besides, the number of parameters (Params), GFLOPs, and FPS are measured to evaluate the scale, computational complexity, and computational efficiency of the model, respectively. Specifically, in the following result tables, $P_{d}$ is expressed in units of $10^{-2}$, $F_{a}$ in $10^{-5}$, Params in 1024 bytes (i.e., KB), GFLOPs in GFLOPs per frame with $480\times 720$ size, and FPS in frames per second.
@@ 4.1.3 Implementation details
The input of our DeepPro was set to 40 consecutive frames ($T=40$) in a sequence, and the number of temporal probes was set to 4 ($m=4$). For data augmentation, we randomly selected one frame as the first frame in the input. If the number of subsequent frames is less than $T$, images with all zero values are filled. We adopted random cropping with a patch size of $128\times 128$, and ensured that the probability of the existence of targets in a sample is higher than 75%. The Soft-IoU [113] loss function was used for loss calculation. In the inference phase, we set the repetition rate at 10% [^5] For example, the first sample is continuous frames from the $1^{st}$ to the $T^{th}$, and the next sample is continuous frames from the $\big{(}(1-10\%)T+1\big{)}^{th}$ to the $\big{(}(1-10\%)T+T\big{)}^{th}$. to ensure that the targets at both ends of the input sequence can be well detected.
@@
Methods $SNR\leq 3$ NUDT-MIRSDT NUDT-MIRSDT-HiNo Params FPS
@@ 4.2 Comparison to the state-of-the-arts
To demonstrate the effectiveness of our method, we compared our method with state-of-the-art IRST detection methods, including traditional MIRST detection methods (MSLSTIPT [6], IMNN-LWEC [74], SRSTT [8], 4DST-BTMD [32], STRL-LBCM [75], 4D-TR [33], 4D-TT [33], and NFTDGSTV [103]), deep learning-based SIRST detection methods (ACM [114], ALCNet [55], Res-UNet [31], DNA-Net [34], ISNet [61], UIUNet [35], AGPCNet [115], MSHNet [116], SCTransNet [36], and RPCANet [37]), and deep learning-based MIRST detection methods (Res-UNet+DTUM [7], STDMANet [26], and Res-UNet+RFR [5]) on the NUDT-MIRSDT dataset and the IRDST-simulation dataset. All traditional methods were implemented with their default parameters. We adopted 0 as the threshold for all deep learning-based methods, i.e., 0.5 after sigmoid operation [62, 7, 117]. Quantitative results are presented in Tables II, III and Fig. 11, and qualitative results are shown in Fig. 12.
@@ 4.2.1 Quantitative results
As shown in Table II and Fig. 11, our DeepPro outperforms the state-of-the-art methods in all scenarios. Especially, on the low-SNR subset and the NUDT-MIRSDT-HiNo dataset, our DeepPro improves $P_{d}$ by $3.02\%$ and $7.52\%$, and reduces $F_{a}$ by $1.15\times 10^{-5}$ and $0.19\times 10^{-5}$ as compared to the second best methods. The results verify that the temporal profile has significant advantages in containing more distinctive characteristics for IRST detection. That is, long-term temporal information is worthy of attention. Besides, among the deep learning-based methods, our DeepPro has the fewest parameters and achieves the best inference efficiency and the lowest GFLOPs per frame. That is because, our DeepPro performs calculations (e.g., addition and multiplication) only in the time dimension for detection, which still benefits from the superiority of the temporal profile.
From the results of the state-of-the-art methods, we can also observe the advantages of temporal information. In most scenarios (e.g., the NUDT-MIRSDT dataset and the NUDT-MIRSDT-HiNo dataset), deep learning-based SIRST detection methods perform better than traditional MIRST detection methods. However, on the low-SNR subset, some traditional methods (e.g., SRSTT, 4D-TR, 4D-TT, and 4DST-BTMD) significantly outperform deep learning-based SIRST detection methods. That is because, it is hard to detect dim targets (e.g., $SNR\leq 3$) without temporal information. Moreover, SRSTT [8] in traditional methods and STDMANet [26] in deep learning-based methods achieve much better detection performance than other existing methods on dim targets and complex scenarios (e.g., the NUDT-MIRSDT-HiNo dataset), because they can extract longer temporal features from more input frames (i.e., 30 and 20, respectively) in a sequence.
Figure 12: Visual comparison on the NUDT-MIRSDT dataset. (a1-a3) Results on the NUDT-MIRSDT ($SNR\leq 3$) subset. (b1-b4) Results on the NUDT-MIRSDT ($SNR>3$) subset. For better visualization, the target area is enlarged in the top-right corner and highlighted with a red circle. The false alarm area is marked with a yellow circle.
@@ 4.2.2 Qualitative results
We compared the qualitative results of different detection methods on the NUDT-MIRSDT dataset, as shown in Fig. 12. Compared with the state-of-the-art methods, our DeepPro generates better visual results, especially on dim targets (e.g., $SNR\leq 3$) due to the influence of temporal profile information. In complex scenarios (e.g., a1, a2, and a3), traditional methods and deep learning-based SIRST methods have significant disadvantages in suppressing false alarms, while most deep learning-based MIRST detection methods (e.g., Res-UNet+DTUM, STDMANet, and our DeepPro) with long sequence as input perform well in this term. That shows the importance of long temporal information for IRST detection.
@@ 4.3.2 Length of temporal profile information
To explore the impact of the length of the temporal profile information used in IRST detection, we changed the number $T$ of input frames to 10, 20, 60, and 80, as well as the size of SCorM (i.e., $T\times T$). The results are summarized in Table V. It can be observed that with the extension of temporal information, $P_{d}$ and $F_{a}$ are significantly improved due to more information for predictions, albeit with a slight increase in computational cost. However, when the length of the time dimension exceeds 40 (e.g., $T=60,80$), there is no significant improvement in the detection performance. Instead, the number of parameters, GFLOPs, and FPS increase significantly.
Figure 14: Visualization of the statistic-based correlation matrixes in the third level with different lengths. (a) SCorMs with length 10 and (b) SCorMs with length 20 are irregular. (c) SCorMs with length 40 and (b) SCorMs with length 60 are approximately symmetric about the minor diagonals. (SCorMs with length 20 are also approximately symmetric about the minor diagonals in the second level as shown in the Appendices. This explains why DeepPro with $T=20$ demonstrates acceptable performance as compared to network with $T=40$, and even performs much better than STDMANet.)
@@
Notably, we compared the results of STDMANet [26] with the input of length 20 in Table II and the results of the variant with $T=20$ in Table V. It can be observed that the detection performance of our DeepPro is obviously superior in all terms of $P_{d}$, $F_{a}$, AUC, Params, and FPS. Meanwhile, the parameters are much smaller, and the computational efficiency is much higher than those in other methods. This demonstrates that the temporal profile information is more important than local spatial-temporal information in IRST detection.

### v2

@@ 3.2.1 Attribution approach for IRST detection networks
$IG(\mathbf{x})=(\mathbf{x}-\mathbf{x^{\prime}})\int_{0}^{1}\frac{\partial F(\mathbf{x^{\prime}}+\alpha\times(\mathbf{x}-\mathbf{x^{\prime}}))}{\partial\mathbf{x}}d\alpha.$ (4)
@@
Based on integrated gradients, we developed the first prediction attribution analysis tool for IRST detection networks. Unlike classification networks, segmentation-based IRST detection networks generate the prediction of each spatial pixel in the input $\mathbf{S}\in\mathbb{R}^{T\times H\times W}$, and can be represented as $D:\mathbb{R}^{T\times H\times W}\longrightarrow\mathbb{R}^{H\times W}$. $T$ denotes the temporal length of the input, while $H\times W$ denotes the input’s spatial size. To interpret $D$, we attribute the prediction of a single target globally. Specifically, we need a converter to quantify the prediction of a target, $C:\mathbb{R}^{H\times W}\longrightarrow\mathbb{R}$. In this work, we achieved the converter by summing the predictive scores of all pixels in this target. That is,
@@
$A_{D,C}^{\gamma}(\mathbf{S},\mathbf{S^{\prime}}):=\int_{0}^{1}\frac{\partial C(D(\gamma(\alpha)))}{\partial\gamma(\alpha)}\times\frac{\partial\gamma(\alpha)}{\partial\alpha}d\alpha,$ (6)
where $\mathbf{S^{\prime}}\in\mathbb{R}^{T\times H\times W}$. Since the targets always present as high-frequency components in the input, the baseline should avoid containing high-frequency components. Meantime, arbitrary fluctuation in the background can affect the target prediction. Therefore, the constant tensor is suitable as a baseline for IRST detection, and the constant should be lower than the gray average of the target since integrated gradients fail to highlight black targets [77]. The path used in this work is a straight line, i.e.,
@@ 3.3 Temporal probe mechanism
Figure 8: Visualization of the structure of TPro. Firstly, temporal probes are employed to pull out complete temporal features of a single pixel from the split input features. Then, multiple learnable SCorMs are applied to the gotten features to extract temporal correlation features. This process is performed pixel-wise across all split input features. Finally, all temporal correlation features are concatenated and fused to generate output features. Due to the learned correlationship among signals, target features are enhanced, and background features are suppressed. Notably, the entire pipeline strictly excludes any space-dimension computations.
Inspired by prior research, we remodel the IRST detection task as a one-dimensional (i.e., time-dimension) signal anomaly detection task, treating target signals as temporal anomalies. As previously analyzed, it is important for the networks to extract long-term temporal correlation information for this task. Ordinary 3D convolution operations focus only on short-term features and require stacking multiple convolution layers to achieve long temporal coverage. However, the temporal receptive field gradually expands as the number of stacked layers increases, and thus a amount of feature extraction computation is still based on short-term features. Moreover, multiple 3D convolutional layers bring substantial computational overhead for large-scale feature tensors. Therefore, in this part, we propose a temporal probe mechanism (TPro) to efficiently extract long-term temporal correlation features and to probe whether the temporal features of a pixel exist anomaly. The structure of TPro is shown in Fig. 8.
@@
As shown in Fig. 8, the input basic features are first divided into $m$ features along the channel dimension, i.e., $F^{b}_{i}\in\mathbb{R}^{\frac{C}{m}\times T\times H\times W}$ (i=1,2,,mi=1,2,...,m). To perform computations on the long-term temporal features of a pixel, temporal probes are utilized to individually extract the complete temporal features of each pixel, i.e., $F^{b}_{i}(x,y)\in\mathbb{R}^{\frac{C}{m}\times T\times 1\times 1}$ ($x\in[1,W],y\in[1,H]$). Subsequently, the correlation matrix is applied to the extracted features for the temporal correlation features $F^{cor}_{i}(x,y)\in\mathbb{R}^{\frac{C}{m}\times T\times 1\times 1}$. The process can be described as
@@
where $W_{i}^{T\times T}$ represents the $i^{th}$ correlation matrix. After repeating the process, the temporal correlation features (i.e., $F^{cor}_{i}$) of all pixels are obtained. To fuse temporal correlation information from different channels and correlation matrixes, a pointwise convolution [11] (P-Conv, i.e., $1\times 1\times 1$ convolution) layer $\phi(\cdot)$ is applied to these features (i.e., $F^{cor}_{1},F^{cor}_{2},...,F^{cor}_{m}$) after concatenation along the channel dimension. This process can be described as
@@ 3.4 Deep temporal probe network for IRST detection
Specifically, tens of frames $\mathbf{S}$ are first fed to a temporal convolution (T-Conv) layer [^4] Each convolution layer consists of a convolution operation, batch normalization, and ReLU activation. with a kernel $\mathbf{W_{T}}\in\mathbb{R}^{l\times 1\times 1}$ ($l=5$). Then, the features are sent to a few identical temporal difference residual blocks (TD-ResBlock) to roughly extract the basic temporal features $F^{b}$. Each TD-ResBlock contains a temporal difference convolution (TD-Conv) layer $\psi(\cdot)$ with a $3\times 1\times 1$ kernel $\mathbf{W_{TD}}$ and a temporal dilation of 2, a T-Conv layer ($l=3$), a P-Conv layer and the skip connection. TD-Conv serves to maintain and emphasize the temporal saliency of targets roughly. Inspired by the temporal-difference design in [94], it accomplishes this by applying convolution to multiple temporal difference features (i.e., $f_{t}-f_{t-1},f_{t}-f_{t+1}]$), which can be described as
@@ 4.1.1 Datasets
TABLE I: Parameters related to noise of the datasets. The standard deviation of noise ($\sigma_{n}$), the distribution range of the multiplicative noise ($\sigma_{g}$), the standard deviation of the additive noise ($\sigma_{o}$), the pattern of non-uniformity, and the average SNR values of the whole dataset and the test set.
Datasets $\sigma_{n}$ $\sigma_{g}$ $\sigma_{o}$ non-uniformity SNR SNR (test)
NUDT-MIRSDT 0.1 0 0 none 11.67 7.00
NUDT-MIRSDT-HiNo 8.0 0.15 1.3 pixel-wise 3.86 3.24
@@ 4.1.2 Metrics
We followed [36] to use the probability of detection ($P_{d}$), false alarm rate ($F_{a}$), and area under curve (AUC) as quantitative metrics for the evaluation of IRST detection performance. $P_{d}$ measures the ability of a method to detect targets. $F_{a}$ measures the ability of a method to correctly identify the background, such as clutter. AUC is the area value under the receiver operating characteristics (ROC) curve [20, 16], reflecting the detection performance under different thresholds. Besides, the number of parameters (#Params), GFLOPs, and FPS are measured to evaluate the scale, computational complexity, and computational efficiency of the model, respectively. Specifically, in the following result tables, $P_{d}$ is expressed in units of $10^{-2}$, $F_{a}$ in $10^{-5}$, #Params in 1024 bytes (i.e., KB), GFLOPs in GFLOPs per frame with $480\times 720$ size, and FPS in frames per second.
@@ 4.1.3 Implementation details
The input of our DeepPro was set to 40 consecutive frames ($T=40$) in a sequence, and the number of temporal probes was set to 4 ($m=4$). For data augmentation, we randomly selected one frame as the first frame in the input. If the number of subsequent frames is less than $T$, images with all zero values are filled. We adopted random cropping with a patch size of $128\times 128$, and ensured that the probability of the existence of targets in a sample is higher than 75%. The Soft-IoU [54] loss function was used for loss calculation. In the inference phase, we set the repetition rate at 10% [^5] For example, the first sample is continuous frames from the $1^{st}$ to the $T^{th}$, and the next sample is continuous frames from the $\big((1-10\%)T+1\big)^{th}$ to the $\big((1-10\%)T+T\big)^{th}$. to ensure that the targets at both ends of the input sequence can be well detected. When the remaining frames are less than $T$, overlapping windows are used to form complete frames. The final detection results for frames input multiple times through overlapping regions are combined using a union operation.
@@
Methods $SNR\leq 3$ NUDT-MIRSDT NUDT-MIRSDT-HiNo #Params FPS
@@ 4.2 Comparison to the state-of-the-arts
To demonstrate the effectiveness of our method, we compared our method with state-of-the-art IRST detection methods, including traditional MIRST detection methods (MSLSTIPT [68], IMNN-LWEC [45], SRSTT [35], 4DST-BTMD [46], STRL-LBCM [47], 4D-TR [81], 4D-$T$ [81], and NFTDGSTV [44]), deep learning-based SIRST detection methods (ACM [13], ALCNet [14], Res-UNet [85], DNA-Net [34], ISNet [104], UIUNet [83], AGPCNet [108], MSHNet [42], SCTransNet [95], and RPCANet [82]), and deep learning-based MIRST detection methods (Res-UNet+DTUM [36], STDMANet [87], and Res-UNet+RFR [91]) on the NUDT-MIRSDT dataset and the IRDST-simulation dataset. All traditional methods were implemented with their default parameters. We adopted 0 as the threshold for all deep learning-based methods, i.e., 0.5 after sigmoid operation [92, 36]. Quantitative results are presented in Tables II, III and Fig. 11, and qualitative results are shown in Fig. 12.
@@ 4.2.1 Quantitative results
As shown in Table II and Fig. 11, our DeepPro outperforms the state-of-the-art methods in all scenarios. Especially, on the low-SNR subset and the NUDT-MIRSDT-HiNo dataset, our DeepPro improves $P_{d}$ by $3.02\%$ and $3.53\%$, and reduces $F_{a}$ by $1.15\times 10^{-5}$ and $0.19\times 10^{-5}$ as compared to the second best results. The results verify that the temporal profile has significant advantages in containing more distinctive characteristics for IRST detection. That is, long-term temporal information is worthy of attention. Besides, among the deep learning-based methods, our DeepPro has the fewest parameters and achieves the best inference efficiency and the lowest GFLOPs per frame. That is because, our DeepPro performs calculations (e.g., addition and multiplication) only in the time dimension for detection, which still benefits from the superiority of the temporal profile.
From the results of the state-of-the-art methods, we can also observe the advantages of temporal information. In most scenarios (e.g., the NUDT-MIRSDT dataset and the NUDT-MIRSDT-HiNo dataset), deep learning-based SIRST detection methods perform better than traditional MIRST detection methods. However, on the low-SNR subset, some traditional methods (e.g., SRSTT, 4D-TR, 4D-$T$, and 4DST-BTMD) significantly outperform deep learning-based SIRST detection methods. That is because, it is hard to detect dim targets (e.g., $SNR\leq 3$) without temporal information. Moreover, SRSTT [35] in traditional methods and STDMANet [87] in deep learning-based methods achieve much better detection performance than other existing methods on dim targets and complex scenarios (e.g., the NUDT-MIRSDT-HiNo dataset), because they can extract longer temporal features from more input frames (i.e., 30 and 20, respectively) in a sequence.
Figure 12: Visual comparison on the NUDT-MIRSDT dataset. (a1-a3) Results on the NUDT-MIRSDT ($SNR\leq 3$) subset. (b1-b3) Results on the NUDT-MIRSDT ($SNR>3$) subset. The first column shows the temporal profile (TP) of the target central pixel. For better visualization, the target area is enlarged in the top-right corner and highlighted with a red circle. The false alarm area is marked with a yellow circle.
@@ 4.2.2 Qualitative results
We compared the qualitative results of different detection methods on the NUDT-MIRSDT dataset, as shown in Fig. 12. Compared with the state-of-the-art methods, DeepPro generates better visual results, especially on dim targets (e.g., $SNR\leq 3$) nearly invisible in the spatial domain, due to the enhanced distinguishability of their signals in the temporal profile. In complex scenarios (e.g., a1, a2, and a3), traditional methods and deep learning-based SIRST methods have significant disadvantages in suppressing false alarms, while most deep learning-based MIRST detection methods (e.g., STDMANet, and our DeepPro) with long sequence as input perform well in this term. That shows the importance of long temporal information for IRST detection.
@@ 4.3.2 Length of temporal profile information
To explore the impact of the length of the temporal profile information used in IRST detection, we changed the number $T$ of input frames to 10, 20, 60, and 80, as well as the size of SCorM (i.e., $T\times T$). The results are summarized in Table V. It can be observed that with the extension of temporal information, $P_{d}$ and $F_{a}$ are significantly improved due to more information for predictions, albeit with a slight increase in computational cost. However, when the length of the time dimension exceeds 40 (e.g., $T=60,80$), there is no significant improvement in the detection performance. Instead, the number of parameters and GFLOPs increase significantly, while FPS decreases noticeably.
Figure 14: Visualization of the statistic-based correlation matrixes in the third level with different lengths. (a) SCorMs with length 10 and (b) SCorMs with length 20 are irregular. (c) SCorMs with length 40 and (b) SCorMs with length 60 are approximately symmetric about the minor diagonals. (SCorMs with length 20 are also approximately symmetric about the minor diagonals in the second level as shown in the Appendices. This explains why DeepPro with $T=20$ demonstrates acceptable performance as compared to the network with $T=40$, and even performs much better than STDMANet.)
@@
Notably, we compared the results of STDMANet [87] with the input of length 20 in Table II and the results of the variant with $T=20$ in Table V. It can be observed that the detection performance of our DeepPro is obviously superior in all terms of $P_{d}$, $F_{a}$, AUC, #Params, and FPS. Meanwhile, the parameters are much smaller, and the computational efficiency is much higher than those in other methods. This demonstrates that the temporal profile information is more important than local spatial-temporal information in IRST detection.